# High-Fidelity SLAM Using Gaussian Splatting with Rendering-Guided Densification and Regularized Optimization

### v1

@@ Abstract
We propose a dense RGBD SLAM system based on 3D Gaussian Splatting that provides metrically accurate pose tracking and visually realistic reconstruction. To this end, we first propose a Gaussian densification strategy based on the rendering loss to map unobserved areas and refine reobserved areas. Second, we introduce extra regularization parameters to alleviate the “forgetting” problem in the continuous mapping problem, where parameters tend to overfit the latest frame and result in decreasing rendering quality for previous frames. Both mapping and tracking are performed with Gaussian parameters by minimizing re-rendering loss in a differentiable way. Compared to recent neural and concurrently developed gaussian splatting RGBD SLAM baselines, our method achieves state-of-the-art results on the synthetic dataset Replica and competitive results on the real-world dataset TUM.
@@ I Introduction
Dense visual SLAM with RGBD inputs is essential for many downstream tasks in mobile robots, AR/VR and robot manipulation. Traditional SLAM systems have focused mainly on camera tracking [1] and geometric surface reconstruction [2]. However, surface appearance reconstruction, which contains rich information for scene understanding, is often lacking in traditional SLAM systems. Motivated by the success of NeRF [3] on novel view synthesis, recent RGBD SLAM methods use volumetric structures [4, 5, 6, 7] or points [8] as the map representation. Combined with neural networks as decoders, these methods can achieve both texture and geometry reconstruction. However, limited by the expensive sampling along the ray in the optimization of NeRF, the above mentioned methods cannot render full-resolution images with high quality (commonly only 0.5% pixels are used for mapping [5], leading to low quality image rendering; see fig. 3 and table I).
The recent work 3D Gaussian Splatting (3DGS) [9] provides a more efficient way for novel view synthesis. Instead of expensive sampling along the ray, 3DGS relies on rasterization for rendering, which accelerates the mapping process so as to handle full-resolution images. The original 3DGS needs prior camera poses, which are often estimated by Structure-from-Motion [10]. In this work, we extend 3DGS in an SLAM framework, removing the need for prior camera poses. Our method processes sequential RGBD frame inputs, simultaneously optimizing Gaussian parameters and estimating camera poses. Though there are some concurrent works combining 3DGS with SLAM [11, 12], we show our method achieves better or competitive results compared to recent baselines.
@@
Our second contribution is the regularized optimization during the mapping. The continuous mapping suffers from the “forgetting” problem, which means that Gaussian parameters tend to overfit the latest frame and lead to decrease of reconstruction quality for previous frames (cf. Figure 2). To alleviate the forgetting problem, we introduce extra parameters to supervise the learning process, and the experiment shows that our method can preserve the rendering quality of previously visited areas.
Experiments show that our method achieves state-of-the-art reconstruction quality and tracking accuracy on Replica dataset and competitive results on the real-world dataset TUM-RGBD. In the ablation study, we show the effectiveness of our method. We also find that continuous mapping can reconstruct high-quality images even if there exists motion blur, which may inspire the 3DGS reconstruction with blurry inputs.
@@ II-A Visual SLAM
Conventional visual SLAM methods are divided into direct and indirect methods. Indirect methods (e.g., ORB-SLAM [1]) first detect points of interest and attach them by feature descriptors. The tracking is conducted by feature matching and minimizing the re-projection error across frames. Direct methods (e.g., DSO [13]) consume raw pixels and construct photometric loss for tracking. These conventional visual SLAM methods focus more on tracking instead of the scene reconstruction. Since NeRF [14] shows powerful scene reconstruction abilities, recent work applies neural representation in the SLAM to produce more complete and photorealistic scene reconstruction. For example, iMAP [7] uses a neural network to represent the scene; NICE-SLAM [4] and later work [5, 6] combine neural networks and voxel grids to improve reconstruction quality. Thanks to the generalization capabilities of neural networks, these neural representations have more powerful abilities to fill holes and construct more complete scenes compared to explicit representations. Although neural representations achieve good results in geometric scene reconstruction, their performance in recovering high-fidelity scenes is not satisfactory enough. In this work, we use parameterized Gaussians [9] as the map representation and achieve a better photorealistic reconstruction. Concurrent with our work, some SLAM systems are also based on Gaussian Splatting: MonoGS [15] expands the map by random sampling; Gaussian-SLAM [16] divides the scene into many submaps; Photo-SLAM [17] conducts tracking by minimizing the reprojection error; SplaTAM [11] and GS-SLAM [12] densify the map only considering the unobserved regions. Our method expands the map by considering filling holes in unobserved areas and refining reobserved areas, which can effectively improve the reconstruction quality. In addition, our regularized optimization can effectively alleviate the forgetting problem during mapping and preserve details of previously visited areas.
@@ II-B Photo-realistic reconstruction
NeRF [14] opens a convenient way for photo-realistic reconstruction from only 2D images. Based on neural networks and volumetric differentiable rendering, NeRF can render novel photorealistic views that were not observed in the inputs. Many follow-up works promote original NeRF by accelerating training speed [18], improving rendering quality [19], etc. Some works [20] also reveal that neural networks are not necessary for novel view synthesis. As opposed to Gaussian Splatting [9], however, all the above-mentioned methods adopt expensive ray-marching when calculating one pixel, resulting in low rendering speed (lower than 15 fps). Gaussian Splatting uses parameterized Gaussians (as introduced in section III) as rendering primitives. Instead of expensive ray-marching, Gaussian Splatting renders the image by tile-based rasterization, which can achieve much faster rendering speed (around 100 fps). The original Gaussian Splatting needs known camera poses (commonly achieved by SfM [10]) and optimizes Gaussian parameters offline. In this work, we consider sequential RGBD frame inputs and optimize parameters online, which releases the requirement of known camera poses and brings the possibility of online applications such as exploration.
@@ III Preliminary: Gaussian Splatting Rendering
We use 3D Gaussians as the map representation primitives for the color and depth image rendering. As described in [9], each Gaussian consists of parameters
@@
where Gaussian means $\bm{\mu}_{i}\in\mathbb{R}^{3}$, scales $\bm{s_{i}}\in\mathbb{R}^{3}$, rotation quaternion $\bm{r}_{i}\in\mathbb{R}^{4}$, RGB color $\bm{c}_{i}\in\mathbb{R}^{3}$, and opacity $o_{i}\in\mathbb{R}$. All of these parameters are optimizable during training.
@@
$C=\sum_{i\in N}\bm{c}_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (2)
$D=\sum_{i\in N}z_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (3)
@@
Though eq. 2 is the same as the volume rendering used in the classical NeRF [14], rasterization-based rendering is much more efficient. Because rasterization allows us to project and render objects in the space directly instead of expensive random sampling along the ray used in a full volumetric representation.
@@ IV Methods
Figure 1 shows an overview of our method. Given RGBD frames and estimated camera poses, we update the map by comparing the rendered images and the ground truth to identify unobserved regions and areas requiring refinement. Regularization terms are incorporated into the optimization process to mitigate the issue of forgetting during mapping (section IV-A). We track the camera pose in the Gaussian map by minimizing color and depth re-rendering loss (section IV-B).
@@ IV-A Mapping
Gaussian Densification. Different from initializing the Gaussian parameters with all images in the original 3D Gaussian Splatting [9], we process sequential RGBD frame inputs. For the latest RGBD input frame with an estimated pose (section IV-B), we need to add new Gaussians to the map, either to complete unobserved regions or to improve the rendering quality of previously mapped regions.
@@
$O=\sum_{i\in N}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (5)
In the opacity image, if the rendered value is below some predefined opacity threshold $\tau_{opa}$ (i.e., $O(x,y)<\tau_{\mathrm{opa}}$), we need to densify the corresponding regions to fill holes.
Second, due to e.g. occlusion and illumination changes, the rendered image is often view-dependent; in addition, the optimized Gaussian parameters may get trapped into local minima to fit certain frames. To account for such view-dependent errors, we propose a kind of rendered error guided densification. When giving the estimated pose, we render the color image $\hat{C}$ and the depth image $\hat{D}$ at this pose; the current reference color image and the depth image are $C$ and $D$. We add new Gaussians when the error between the rendered image and the input image is large, where the color error and depth error are $E_{\mathrm{color}}(x,y)=|\hat{C}(x,y)-C(x,y)|$ and $E_{\mathrm{depth}}(x,y)=|\hat{D}(x,y)-D(x,y)|$ respectively. We will densify those regions where $E_{\mathrm{color}}(x,y)>\tau_{\mathrm{color}}$ or $\frac{E_{\mathrm{depth}}(x,y)}{D(x,y)}>\tau_{\mathrm{depth}}$, where $\tau_{\mathrm{color}}$ and $\tau_{\mathrm{depth}}$ are predefined thresholds.
In summary, we densify the regions where $\big{(}O(x,y)<\tau_{\mathrm{opa}}\big{)}\ |\ \big{(}E_{\mathrm{color}}(x,y)>\tau_{\mathrm{color}}\big{)}\ |\ \big{(}\frac{E_{\mathrm{depth}}(x,y)}{D(x,y)}>\tau_{\mathrm{depth}}\big{)}$. With depth images, we project corresponding depths to the space directly. The experiment (table IV) shows our densification strategy can significantly improve the map quality.
Continuous Mapping. We optimize Gaussian parameters by minimizing the loss $\ell_{1}$ between the rendered color and depth image and the input frames; in addition, following the original implementation in [9], we add an extra $\mathrm{SSIM}$ term.
@@
As a continuous learning problem, the incremental mapping in our work also suffers from the forgetting problem [21]. That is, the same Gaussian might be “seen” in multiple camera poses, but the Gaussian parameters tend to be optimized (overfitted) to fit the latest frame, resulting in a performance drop for previous frames (as demonstrated in fig. 2).
@@
To alleviate the overfitting problem, prior work [4, 5, 6] maintains a keyframe buffer. When optimizing the latest frame, some previous keyframes are selected from the buffer based on some criteria such as overlap ratio with the current frame. The selected keyframes will then be optimized together with the current frame. In the original 3D Gaussian Splatting, all frames are queued and picked randomly for many iterations to optimize Gaussian parameters.
We can, of course, pick many previous frames whenever we optimize the current frame; however, optimizing with multiple frames will lead to more time consumption, which is not suitable for online continual mapping. To prevent Gaussian parameters from forgetting, we add an extra regularization term $\mathcal{L}_{\mathrm{reg}}$ when mapping:
@@
where $G$ refers to the set of Gaussians involved in the optimization for the current frame; $\bm{s}_{i}^{t},\bm{r}_{i}^{t},z_{i}^{t}$ are scales, colors and depth to be optimized; $\bm{s}_{i}^{*},\bm{r}_{i}^{*},z_{i}^{*}$ are current parameter values. $\Omega$’s are importance weights for different parameters.
To calculate importance weights, for each Gaussian represented by Equation 1, we additionally add the following parameters:
@@
where $N_{i}^{\mathrm{seen}}$ records the “seen times” of each Gaussian; $\Sigma$ are the sum of gradients of $\mathcal{L}_{\mathrm{mapping}}$ with respect to parameters, representing the sensitivity to changes. After training each frame, we update the importance weights by averaging the gradients of mapping loss with respect to the parameters. Taking the scale s for example:
$\Omega_{i}^{\bm{s}}=\frac{\Sigma_{i}^{\bm{s}}}{N_{i}^{\mathrm{seen}}}=\frac{1}{N_{i}^{\mathrm{seen}}}\sum_{i}^{N_{i}^{\mathrm{seen}}}(\frac{\partial\mathcal{L}_{\mathrm{mapping}}}{\partial\bm{s}})$ (9)
@@
$\mathcal{L}_{\mathrm{{mapping}}}^{\prime}=\mathcal{L}_{\mathrm{mapping}}+\mathcal{L}_{\mathrm{reg}}$ (10)
@@ IV-B Tracking
We conduct camera tracking in a render-and-compare way similar to that of iNeRF [3] for every input frame. The camera pose is parameterized by $T=(t,q)$, including the 3D translation $t$ and the rotation quaternion $q$. When tracking the new frame, we initialize the camera pose with the assumption of constant camera velocity. Since the color image can be affected by light, we decouple lightness from the image during tracking, similar to [22]. Specifically, let $\rho$ denote the operation of converting the image from the RGB space to the LAB space, and we discard the light channel (L channel) when comparing the two images.

### v2

@@ Abstract
We propose a dense RGBD SLAM system based on 3D Gaussian Splatting that provides metrically accurate pose tracking and visually realistic reconstruction. To this end, we first propose a Gaussian densification strategy based on the rendering loss to map unobserved areas and refine reobserved areas. Second, we introduce extra regularization parameters to alleviate the “forgetting” problem during contiunous mapping, where parameters tend to overfit the latest frame and result in decreasing rendering quality for previous frames. Both mapping and tracking are performed with Gaussian parameters by minimizing re-rendering loss in a differentiable way. Compared to recent neural and concurrently developed Gaussian splatting RGBD SLAM baselines, our method achieves state-of-the-art results on the synthetic dataset Replica and competitive results on the real-world dataset TUM. The code is released on https://github.com/ljjTYJR/HF-SLAM.
@@ I Introduction
Dense visual SLAM with RGBD inputs is essential for many downstream tasks in mobile robots, AR/VR and robot manipulation. Traditional SLAM systems have focused mainly on camera tracking [mur2015orb] and geometric surface reconstruction [newcombe2011kinectfusion]. However, surface appearance reconstruction, which contains rich information for scene understanding, is often lacking in traditional SLAM systems. Motivated by the success of NeRF [yen2021inerf] on novel view synthesis, recent RGBD SLAM methods use volumetric structures [zhu2022nice, johari2023eslam, wang2023co, sucar2021imap] or points [sandstrom2023point] as the map representation. Combined with neural networks as decoders, these methods can achieve both appearance and geometry reconstruction. However, limited by the computationally expensive sampling along rays in the optimization of NeRF, the above mentioned methods cannot render full-resolution images with high quality (commonly only 0.5% pixels are used for mapping [johari2023eslam], leading to low quality image rendering; see Fig. 3 and Tab. I).
The recent work 3D Gaussian Splatting (3DGS) [kerbl20233dgaussian] provides a more efficient way for novel view synthesis. Instead of expensive sampling along the ray, 3DGS relies on rasterization for rendering, which accelerates the mapping process so as to handle full-resolution images. The original 3DGS needs prior camera poses, which are often estimated by Structure-from-Motion [schonberger2016structure]. In this work, we extend 3DGS to the case of online tracking and mapping, removing the need for prior camera poses. Our method processes sequential RGBD frame inputs, simultaneously optimizing Gaussian parameters and estimating camera poses. Though there are some concurrent works combining 3DGS with SLAM [keetha2023splatam, yan2023gs], we show our method achieves better or competitive rendering and tracking results compared to recent baselines.
@@
Our second contribution is regularized optimization during mapping. Continuous mapping with 3DGS suffers from the “forgetting” problem, which means that Gaussian parameters tend to overfit to the latest frame and lead to decreased reconstruction quality for previous frames (see Fig. 2). To alleviate the forgetting problem, we introduce extra parameters to supervise the learning process, and our experiments show that our method can preserve the rendering quality of previously visited areas.
We demonstrate state-of-the-art reconstruction quality and tracking accuracy on Replica dataset and competitive results on the real-world dataset TUM-RGBD. We show the relative effectiveness of the two main contributions in the ablation study. We also find that continuous mapping can reconstruct high-quality images even if when faced with motion blur, which may inspire work on 3DGS reconstruction with blurry inputs.
@@ II-A Visual SLAM
Conventional visual SLAM methods are divided into direct and indirect methods. Indirect methods (e.g., ORB-SLAM [mur2015orb]) first detect points of interest and attach feature descriptors. Tracking is conducted by feature matching and minimizing the re-projection error across frames. Direct methods (e.g., DSO [engel2017direct]) consume raw pixels and construct photometric loss for tracking. These conventional visual SLAM methods focus more on tracking instead of scene reconstruction. Since NeRF [mildenhall2021nerf] has shown powerful scene reconstruction abilities, recent work applies neural representation in SLAM to produce more complete and photorealistic scene reconstruction. For example, iMAP [sucar2021imap] uses a neural network to represent the scene; NICE-SLAM [zhu2022nice] and later work [johari2023eslam, wang2023co] combine neural networks and voxel grids to improve reconstruction quality. Thanks to the generalization capabilities of neural networks, these neural representations have more powerful abilities to fill holes and construct more complete scenes compared to explicit representations. Although neural representations achieve good results in geometric scene reconstruction, their performance in recovering high-fidelity scenes is not satisfactory enough. In this work, we use parameterized Gaussians [kerbl20233dgaussian] as the map representation and achieve a more photorealistic reconstruction. Concurrent with our work, some SLAM systems are also based on Gaussian Splatting: MonoGS [matsuki2023gaussian] expands the map by random sampling; Gaussian-SLAM [yugay2023gaussian] divides the scene into many submaps; Photo-SLAM [huang2023photo] conducts tracking by minimizing the reprojection error; SplaTAM [keetha2023splatam] and GS-SLAM [yan2023gs] densify the map only considering the unobserved regions (more information can be found in [chen2024survey]). Our method expands the map by considering filling holes in unobserved areas and refining reobserved areas, which can effectively improve the reconstruction quality. In addition, our regularized optimization can effectively alleviate the forgetting problem during mapping and preserve details of previously visited areas.
@@ II-B Photo-realistic reconstruction
NeRF [mildenhall2021nerf] opens a convenient way for photo-realistic reconstruction from only 2D images. Based on neural networks and volumetric differentiable rendering, NeRF can render novel photorealistic views that were not observed in the inputs. Many follow-up works improve on the original NeRF by accelerating training speed [muller2022instant], improving rendering quality [barron2023zip], etc. Some works [fridovich2022plenoxels] also reveal that neural networks are not necessary for novel view synthesis. As opposed to Gaussian splatting [kerbl20233dgaussian], however, all the above-mentioned methods adopt expensive ray-marching when calculating one pixel, resulting in low rendering speed (lower than 15 fps). Gaussian Splatting uses parameterized Gaussians (as introduced in Sec. III) as rendering primitives. Instead of expensive ray-marching, Gaussian Splatting renders the image by tile-based rasterization, which can achieve much faster rendering speed (around 100 fps). The original Gaussian Splatting needs known camera poses (commonly achieved by SfM [schonberger2016structure]) and optimizes Gaussian parameters offline. In this work, we consider sequential RGBD frame inputs and optimize parameters online, which releases the requirement of known camera poses and brings the possibility of online applications such as exploration.
@@ III Preliminary: Gaussian Splatting Rendering
We use 3D Gaussians as the map representation primitives for the color and depth image rendering. As described in [kerbl20233dgaussian], each Gaussian consists of parameters
@@
with Gaussian means $\bm{\mu}_{i}\in\mathbb{R}^{3}$, scales $\bm{s_{i}}\in\mathbb{R}^{3}$, rotation quaternion $\bm{r}_{i}\in\mathbb{R}^{4}$, RGB color $\bm{c}_{i}\in\mathbb{R}^{3}$, and opacity $o_{i}\in\mathbb{R}$. All of these parameters are optimizable during training.
@@
$C=\sum_{i\in N}\bm{c}_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (2)
$D=\sum_{i\in N}z_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (3)
@@
Though Eq. 2 is the same as the volume rendering used in the classical NeRF [mildenhall2021nerf], rasterization-based rendering is much more efficient. Because rasterization allows us to project and render objects in the space directly instead of expensive sampling along the ray used in a full volumetric representation.
@@ IV Methods
Figure 1 shows an overview of our method. Given RGBD frames and estimated camera poses, we update the map by comparing the rendered images and the ground truth to identify unobserved regions and areas requiring refinement. Regularization terms are incorporated into the optimization process to mitigate the issue of forgetting during mapping (Sec. IV-A). We track the camera pose in the Gaussian map by minimizing color and depth re-rendering loss (Sec. IV-B).
@@ IV-A Mapping
Gaussian Densification. Differently from initializing the Gaussian parameters with all images as in the original 3D Gaussian Splatting [kerbl20233dgaussian], we process sequential RGBD frame inputs. For the latest RGBD input frame with an estimated pose (Sec. IV-B), we need to add new Gaussians to the map, either to complete unobserved regions or to improve the rendering quality of previously mapped regions.
@@
$O=\sum_{i\in N}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}).$ (5)
In the opacity image, if the rendered value is below some predefined opacity threshold $\tau_{\mathrm{opa}}$ (i.e., $O(x,y)<\tau_{\mathrm{opa}}$), we need to densify the corresponding regions to fill holes.
Second, due to e.g. occlusion and illumination changes, the rendered image is often view-dependent. In addition, the optimized Gaussian parameters may get trapped into local minima to fit certain frames. To account for such view-dependent errors, we propose a densification step guided by rendering error. When giving the estimated pose, we render the color image $\hat{C}$ and the depth image $\hat{D}$ at this pose; the current reference color image and the depth image are $C$ and $D$. We add new Gaussians when the error between the rendered image and the input image is large, where the color error and depth error are $E_{\mathrm{color}}(x,y)=|\hat{C}(x,y)-C(x,y)|$ and $E_{\mathrm{depth}}(x,y)=|\hat{D}(x,y)-D(x,y)|$ respectively. We will densify those regions where $E_{\mathrm{color}}(x,y)>\tau_{\mathrm{color}}$ or $\frac{E_{\mathrm{depth}}(x,y)}{D(x,y)}>\tau_{\mathrm{depth}}$, where $\tau_{\mathrm{color}}$ and $\tau_{\mathrm{depth}}$ are predefined thresholds—we divide by $D(x,y)$ when comparing to $\tau_{\text{depth}}$ to compensate for varying scales.
In summary, we densify the regions where $\big{(}O(x,y)<\tau_{\mathrm{opa}}\big{)}\ |\ \big{(}E_{\mathrm{color}}(x,y)>\tau_{\mathrm{color}}\big{)}\ |\ \big{(}\frac{E_{\mathrm{depth}}(x,y)}{D(x,y)}>\tau_{\mathrm{depth}}\big{)}$. With depth images, we project corresponding depths directly. The experiment reported in Tab. IV shows that our densification strategy can significantly improve map quality.
[…]
@@
As a continuous learning problem, the incremental mapping in our work also suffers from the forgetting problem [mccloskey1989catastrophic]: the same Gaussian might be “seen” in multiple camera poses, and the Gaussian parameters tend to be optimized (overfitted) to fit the latest frame, resulting in a performance drop for previous frames (see Fig. 2).
@@
To alleviate the overfitting problem, prior work [zhu2022nice, johari2023eslam, wang2023co] maintains a keyframe buffer from which previous keyframes are selected (based on criteria such as overlap ratio with the current frame) and the selected keyframes are then optimized together with the current frame. In the original 3D Gaussian Splatting, all frames are queued and picked randomly for many iterations to optimize Gaussian parameters.
We can, of course, pick many previous frames whenever we optimize the current frame; however, optimizing with multiple frames leads to longer processing times, which is not suitable for online continual mapping. Instead, to prevent Gaussian parameters from forgetting, we add an extra regularization term $\mathcal{L}_{\mathrm{reg}}$ when mapping:
@@
where $G$ refers to the set of Gaussians involved in the optimization for the current frame; $\bm{s}_{i}^{t},\bm{r}_{i}^{t},z_{i}^{t}$ are scales, colors and depth to be optimized at the step $t$; $\bm{s}_{i}^{*},\bm{r}_{i}^{*},z_{i}^{*}$ are current parameter values, that is, values achieved after step $t-1$. $\Omega$’s are importance weights for different parameters. In this way, we only need to maintain a small keyframe buffer but achieve better reconstruction quality.
To calculate importance weights, for each Gaussian represented by Eq. 1, we additionally add the following parameters:
@@
where $N_{i}^{\mathrm{seen}}$ records the “seen times” of each Gaussian; $\Sigma$ are the sum of gradients of $\mathcal{L}_{\mathrm{mapping}}$ with respect to parameters, representing the sensitivity to changes. After training each frame, we update the importance weights by averaging the gradients of mapping loss with respect to the parameters. Taking the $i$th scale parameter $\textbf{s}_{i}$ as an example, after training with $N_{i}^{\mathrm{seen}}$ frames, its importance weight $\Omega_{i}^{\bm{s}}$ is:
$\Omega_{i}^{\bm{s}}=\frac{\Sigma_{i}^{\bm{s}}}{N_{i}^{\mathrm{seen}}}=\frac{1}{N_{i}^{\mathrm{seen}}}\sum_{j}^{N_{i}^{\mathrm{seen}}}(\frac{\partial\mathcal{L}_{\mathrm{mapping}}}{\partial\bm{s}_{i}})_{j}$ (9)
@@
$\mathcal{L}_{\mathrm{{mapping}}}^{\prime}=\mathcal{L}_{\mathrm{mapping}}+\mathcal{L}_{\mathrm{reg}}.$ (10)
@@ IV-B Tracking
We conduct camera tracking in a render-and-compare way similar to that of iNeRF [yen2021inerf] for every input frame. The camera pose is parameterized by $T=(t,q)$, including the 3D translation $t$ and the rotation quaternion $q$. When tracking the new frame, we initialize the camera pose assuming constant camera velocity. Since the color image can be affected by light, we decouple lightness from the image during tracking, similar to [wang2020self6d]. Specifically, let $\rho$ denote the operation of converting the image from the RGB space to the LAB space, and we discard the light channel (L channel) when comparing the two images.